Let u=(2,6,-7), v=(-1,-1,8), and 3. If (2, 14, 11)=ku+lv what is the value of l?


1
Expert's answer
2021-08-10T09:23:21-0400

Given,

"u=(2,6-7) \space \&\space v=(-1,-1,8)"

"\& k=3"

"Also,\space (2,14,11)=ku+lv"

"=3(2,6,-7)+l(-1,-1,8)"

"(2,14,11)=(6-l,18-l,-21+8l)" "[l(a,b)k(c,d)=(la+kc,lb+kd)]"

"6-l=2 \implies l=4,\\\space also \space 18-l=14\implies l=4\\also\space -21+8l=11\implies l=4\\\therefore l=4"